# Transformers for Stratified Spectropolarimetric Inversion: Proof of Concept

### v2

@@ Abstract
Solar spectropolarimetric inversion—inferring atmospheric conditions from the Stokes vector—is a key diagnostic tool for understanding solar magnetism, but traditional inversion methods are computationally expensive and sensitive to local minima. Advances in artificial intelligence (AI) offer faster solutions, but are often restricted to shallow models or a few spectral lines. We present a proof-of-concept study using a transformer machine learning (ML) model for multi-line, full-Stokes inversion, to infer stratified parameters from synthetic spectra produced from 3D magnetohydrodynamic simulations. We synthesise a large set of Stokes vectors using forward modelling across 15 spectral lines spanning the deep photosphere towards the chromosphere. The model maps full-Stokes input to temperature, magnetic field strength, inclination, azimuth (encoded as $\sin 2\phi$, $\cos 2\phi$), and line-of-sight velocity as a function of optical depth. The transformer incorporates an attention mechanism that allows the model to focus on the most informative regions of the spectrum for each inferred parameter, and uses positional embedding to encode wavelength and depth order. We benchmark it against a multilayer perceptron (MLP), test robustness to noise, and assess generalisation. The transformer outperforms the MLP, especially in the higher layers and for magnetic parameters, yielding higher correlations and more regularised stratifications. The model retains strong performance across a range of noise levels typical for real observations, with magnetic parameter inference degrading predictably while temperature and velocity remain stable. We establish transformer architectures as a powerful tool for spectropolarimetric inversion. This approach paves the way for analysis of large datasets from large solar telescopes.
@@
software: PyTorch (A. Paszke et al., 2019), Matplotlib (J. D. Hunter, 2007), Numpy (C. R. Harris et al., 2020), Astropy (Astropy Collaboration et al., 2022).
@@ 1 Introduction
Inverting the Stokes vector to retrieve physical conditions in the solar atmosphere is fundamental for understanding solar magnetism (J. C. del Toro Iniesta & B. Ruiz Cobo, 2016). Inversion is the inference of depth-dependent atmospheric parameters from the wavelength dependence of the polarized spectrum in spectral lines and corresponding continua. While traditional inversion methods based on radiative transfer are physically grounded, they are also computationally intensive and sensitive to initialisation and noise (A. Asensio Ramos et al., 2012). Crucially, the inversion problem is well suited to machine learning (ML): it can be framed as a supervised regression task with known forward physics, allowing large synthetic training sets with perfect labels, and a deterministic mapping from input (Stokes vectors as a function of wavelength) to output (atmospheric parameters as a function of geometric height or optical depth). These properties make spectropolarimetric inversion an ideal application for ML architectures, especially those that can handle high-dimensional inputs and uncertain mappings.
ML has been widely applied in solar physics, particularly to speed up the retrieval of atmospheric parameters from observed Stokes vectors (A. Asensio Ramos & C. J. D´ıaz Baso, 2019; A. Asensio Ramos et al., 2023). A common approach involves using inversions of a small subset of Stokes vectors and associated atmospheric models as training data for a Multi-Layer Perceptron (MLP) or Convolutional Neural Network (CNN), which can then rapidly provide atmospheric models for the rest of the dataset (I. Milić & R. Gafeira, 2020). However, a major challenge is providing a diverse enough training set to achieve sufficient generalisation, necessitating large data volumes (K. E. Yang et al., 2024). R. Gafeira et al. (2021) improved the accuracy of their inversions by using models produced from CNNs as initial conditions for traditional inversions. H. Socas-Navarro & A. Asensio Ramos (2021) used only forward synthesis, an MLP, and randomly generated atmospheric models to infer target parameters, demonstrating generalisation to unseen Hinode data. Another recent advance is the use of neural networks influenced by physical equations, such as the approach taken by R. Jarolim et al. (2025), C. U. Keller (2025), or C. J. D´ıaz Baso et al. (2025).
Transformers, originally developed for natural language processing (A. Vaswani et al., 2017), have demonstrated success across many domains due to their ability to model long-range dependencies and flexible input structures. By applying a transformer-based architecture to the spectropolarimetric inversion problem, we aim to capture subtle spectral features and inter-wavelength correlations.
@@ 1.1 Aims and hypothesis
The move from MLPs or CNNs to transformers fundamentally changes what the model can learn. Transformers allow the model to attend to non-local spectral features across the ordered line profile, which is especially important in spectropolarimetry where physical information (like Zeeman splitting or velocity gradients) can manifest in subtle, non-local ways. By “non-local”, we refer to correlations or informative relationships between widely separated wavelength points. Unlike MLPs, which treat each wavelength as an independent input feature and do not have awareness of their order, transformers can dynamically learn the spectral context. By “context”, we mean features whose interpretation depends on the broader spectral profiles. To our knowledge, at the time of submission, this was the first application of an attention-based (transformer) neural network to the problem of spectropolarimetric inversion in solar physics. Shortly following submission, A. Asensio Ramos & J. de la Cruz Rodriguez (2025) applied a transformer architecture to the inversion problem with a neural model used to compress the data into a lower dimensionality. Our central hypothesis is that a transformer-based architecture will achieve more accurate spectropolarimetric inversions than an MLP baseline, due to its ability to model spectral order (via learned positional embeddings) and non-local spectral dependencies (via attention mechanisms). For our study, we choose to benchmark the transformer to an MLP because without an attention mechanism the transformer architecture reduces, albeit with some caveats, to a very deep stack of MLPs (i.e. feed-forward networks). This allows us to isolate the impact of the attention mechanism and positional embeddings on inversion performance most effectively and makes a well-tuned MLP the most natural baseline model.
@@ 2.1 Simulation snapshots
For this study, we employ three-dimensional snapshots generated with the MANCHA code (E. Khomenko et al., 2017), as described by M. Modestov et al. (2024). By choosing a quiet Sun simulation, we deliberately select one of the most challenging regimes for spectropolarimetric inversion, where polarisation signals, particularly in the chromosphere, are weak. The simulation starts from an initially field-free state, allowing the Biermann battery effect and a local turbulent dynamo to self-consistently generate magnetic fields reaching strengths comparable to those observed in the quiet Sun. These simulation snapshots have also been used in previous works such as C. Quintero Noda et al. (2023); R. J. Campbell et al. (2025). The computational domain covers an area of $11.52\times 11.52$ Mm2 (resolved with $1152\times 1152$ pixels2), and extends vertically from $1$ Mm beneath the photosphere, with a $1400$ km upper limit. The grid spacing is $10$ km in the horizontal directions and $7$ km vertically. The mean surface magnetic field strength in the saturated regime is approximately $100$ G.
@@ 2.2 Forward synthesis
To generate the synthetic observables for this study, we use the DeSIRe code (B. Ruiz Cobo et al., 2022) to forward-synthesise a set of spectral lines covering atmospheric heights from the deep photosphere to the chromosphere, enabling a comprehensive assessment of model performance at different optical depths.
The simulation outputs are provided on a three-dimensional grid in geometrical height (km). To facilitate forward synthesis with the DeSIRe code, which requires an optical depth scale, we converted the atmospheric stratifications to the required $\log(\tau)$ grid. This was done using the modelador.x utility distributed with DeSIRe, following the methodology outlined in R. J. Campbell et al. (2021). The final grid has $81$ depth points.
Figure 1: Mean Stokes $I$ profile across the full field of view, averaged over all spatial pixels in the MANCHA simulation snapshot used for training. The vertical dashed lines and labels indicate the most important diagnostic spectral lines. These lines span a broad range of atmospheric heights, from the deep photosphere to the mid-chromosphere, providing constraints for inversion at multiple depths.
@@ 3.2 Transformer architecture
In the transformer framework, the core computation within each attention layer involves three components: queries, keys, and values, as originally introduced by A. Vaswani et al. (2017). For a given position in the sequence (e.g., a specific wavelength), the query vector determines what the model is looking for, the key vectors determine where that information is located in the input, and the value vectors represent what is retrieved. For a given input sequence $X\in\mathbb{R}^{N_{\lambda}\times d}$ after projection, the queries, keys, and values are computed as
@@ 3.4 Training and validation
where $\hat{y}_{i}$ and $y_{i}$ are the predicted and true outputs for the $i$-th training example, and $N$ is the number of examples in the batch. We experimented with both the Adam and AdamW optimizers (D. P. Kingma & J. Ba, 2017). The latter includes a modest weight decay ($1\times 10^{-4}$) to gently penalize overly large weights and improve generalization. We used mixed-precision (16-bit) arithmetic with automatic loss scaling, which speeds up training and reduces memory without sacrificing stability. To prevent any single batch from causing a sudden, oversized update, we cap the combined size of all gradient values at $1.0$ so if the raw gradients exceed this threshold, they’re uniformly scaled down such that the model only makes a controlled adjustment. We performed early stopping based on validation loss, with a maximum of 35 epochs per run.
@@ 5 Discussion
In this proof of concept study, we have demonstrated the application of a transformer-based neural network, originally developed for natural language processing, to the inversion problem of inferring stratified atmospheric parameters from full-Stokes, multi-line spectropolarimetric data generated from realistic 3D MHD simulations. Our results show that the transformer model outperforms an MLP baseline in nearly all aspects: reproducing the spatial and vertical morphology of temperature, magnetic field strength, and velocity; achieving higher correlations with the ground truth at all optical depths; and, crucially, delivering physically plausible solutions even in challenging chromospheric layers.
@@
The attention mechanism at the core of the transformer architecture enables the model to leverage non-local spectral features, potentially corresponding to Zeeman splitting or subtle asymmetries, which may be overlooked by models that treat each wavelength as just another input feature, with no explicit bias toward local or long-range spectral relationships. Although a detailed interpretability study is beyond the current scope, initial inspection of the attention maps suggests that the model prioritises physically meaningful regions of the spectrum (especially the most magnetically sensitive Fe I lines) for inference. This could open new avenues for understanding how neural networks extract physical parameters from complex solar spectra.
@@ 6 Outlook
This study provides the first transformer-based model for full-vector Stokes inversions in solar physics, enabling accurate and efficient inference of atmospheric parameters over many spectral lines and at many optical depths. We anticipate a follow-up study to apply the transformer to real spectropolarimetric observations, such as those from Sunrise III, DKIST, and other current or next-generation facilities. Training on simulations with advanced chromospheric physics will be critical to ensure robust generalisation. We may find that the transformer is most useful in providing a close initial guess for subsequent inversion.
R.J.C. would like to thank Robert Ryans and Ernst de Mooij for approving use of an NVIDIA Quadro RTX 8000 GPU for model development, as well as technical staff from the Kelvin2 HPC support teams in using their NVIDIA H100 GPUs. R.J.C. thanks Elena Khomenko for permission to use the MANCHA models. This is UKRI-supported research under the Science and Technology Facilities Council (STFC) grants ST/P000304/1 and ST/X000923/1. Data availability statement: The code that defined the models generated in this study is available on GitHub. The MANCHA simulation code is available on their public repository. The simulation data used in this study were generated by a third party and are not publicly archived. Access to these data may be available from the original authors on request. {contribution} R.J.C. conceived the project, designed and implemented the machine learning models, conducted the analysis and authored the paper. $M.$ provided critical feedback on the framing of the study and the manuscript. C.Q.N. provided simulation data converted to an optical depth scale that formed the training and test set. All authors approved the manuscript before submission.

### v3

@@ Abstract
Solar spectropolarimetric inversion—inferring atmospheric conditions from the Stokes vector—is a key diagnostic tool for understanding solar magnetism, but traditional inversion methods are computationally expensive and sensitive to local minima. Advances in artificial intelligence (AI) offer faster solutions, but are often restricted to shallow models or a few spectral lines. We present a proof-of-concept study using a transformer machine learning (ML) model for multi-line, full-Stokes inversion, to infer stratified parameters from synthetic spectra produced from 3D magnetohydrodynamic simulations. We synthesise a large set of Stokes vectors using forward modelling across 15 spectral lines spanning the deep photosphere towards the chromosphere. The model maps full-Stokes input to temperature, magnetic field strength, inclination, azimuth (encoded as $\sin 2\phi$, $\cos 2\phi$), and line-of-sight velocity as a function of optical depth. The transformer incorporates an attention mechanism that allows the model to focus on the most informative regions of the spectrum for each inferred parameter, and uses positional embedding to encode wavelength and depth order. We benchmark it against a multilayer perceptron (MLP), test robustness to noise, and assess generalisation. The transformer outperforms the MLP, especially in the higher layers and for magnetic parameters, yielding higher correlations and more regularised stratifications. The model retains strong performance across a range of noise levels typical for real observations, with magnetic parameter inference degrading predictably while temperature and velocity remain stable. We explore attention maps, linking the transformer’s learned behaviour to line-formation physics.
@@
software: PyTorch (pytorch2019), Matplotlib (mpl), Numpy (numpy), Astropy (astropy3).
@@ 1 Introduction
Inverting the Stokes vector to retrieve physical conditions in the solar atmosphere is fundamental for understanding solar magnetism (inversions2016). Inversion is the inference of depth-dependent atmospheric parameters from the wavelength dependence of the polarized spectrum in spectral lines and corresponding continua. While traditional inversion methods based on radiative transfer are physically grounded, they are also computationally intensive and sensitive to initialisation and noise (ramos2012). Crucially, the inversion problem is well suited to machine learning (ML): it can be framed as a supervised regression task with known forward physics, allowing large synthetic training sets with perfect labels, and a deterministic mapping from input (Stokes vectors as a function of wavelength) to output (atmospheric parameters as a function of geometric height or optical depth). These properties make spectropolarimetric inversion an ideal application for ML architectures, especially those that can handle high-dimensional inputs and uncertain mappings.
ML has been widely applied in solar physics, particularly to speed up the retrieval of atmospheric parameters from observed Stokes vectors (Ramos2019; RamosLRSP). A common approach involves using inversions of a small subset of Stokes vectors and associated atmospheric models as training data for a Multi-Layer Perceptron (MLP) or Convolutional Neural Network (CNN), which can then rapidly provide atmospheric models for the rest of the dataset (Milic2020). However, a major challenge is providing a diverse enough training set to achieve sufficient generalisation, necessitating large data volumes (SPIn4D). Gafeira2021 improved the accuracy of their inversions by using models produced from CNNs as initial conditions for traditional inversions. Navarro used only forward synthesis, an MLP, and randomly generated atmospheric models to infer target parameters, demonstrating generalisation to unseen Hinode data. Another recent advance is the use of neural networks influenced by physical equations, such as the approach taken by PINNME, keller, or Ramos2025.
Transformers, originally developed for natural language processing (vaswani2017attention), have demonstrated success across many domains due to their ability to model long-range dependencies and flexible input structures. By applying a transformer-based architecture to the spectropolarimetric inversion problem, we aim to capture subtle spectral features and inter-wavelength correlations.
@@ 1.1 Aims and hypothesis
The move from MLPs or CNNs to transformers fundamentally changes what the model can learn. Transformers allow the model to attend to non-local spectral features across the ordered line profile, which is especially important in spectropolarimetry where physical information (like Zeeman splitting or velocity gradients) can manifest in subtle, non-local ways. By “non-local”, we refer to correlations or informative relationships between widely separated wavelength points. Unlike MLPs, which treat each wavelength as an independent input feature and do not have awareness of their order, transformers can dynamically learn the spectral context. By “context”, we mean features whose interpretation depends on the broader spectral profiles. To our knowledge, at the time of submission, this was the first application of an attention-based (transformer) neural network to the problem of spectropolarimetric inversion in solar physics. Shortly following submission, ramos_jaime_2025 applied a transformer architecture to the inversion problem with a neural model used to compress the data into a lower dimensionality. Our central hypothesis is that a transformer-based architecture will achieve more accurate spectropolarimetric inversions than an MLP baseline, due to its ability to model spectral order (via learned positional embeddings) and non-local spectral dependencies (via attention mechanisms). For our study, we choose to benchmark the transformer to an MLP because without an attention mechanism the transformer architecture reduces, albeit with some caveats, to a very deep stack of MLPs (i.e. feed-forward networks). This allows us to isolate the impact of the attention mechanism and positional embeddings on inversion performance most effectively and makes a well-tuned MLP the most natural baseline model.
@@ 2.1 Simulation snapshots
For this study, we employ three-dimensional snapshots generated with the MANCHA code (MANCHA1), as described by MANCHA2. By choosing a quiet Sun simulation, we deliberately select one of the most challenging regimes for spectropolarimetric inversion, where polarisation signals, particularly in the chromosphere, are weak. The simulation starts from an initially field-free state, allowing the Biermann battery effect and a local turbulent dynamo to self-consistently generate magnetic fields reaching strengths comparable to those observed in the quiet Sun. These simulation snapshots have also been used in previous works such as Carlos2023; Campbell2025. The computational domain covers an area of $11.52\times 11.52$ Mm2 (resolved with $1152\times 1152$ pixels2), and extends vertically from $1$ Mm beneath the photosphere, with a $1400$ km upper limit. The grid spacing is $10$ km in the horizontal directions and $7$ km vertically. The mean surface magnetic field strength in the saturated regime is approximately $100$ G.
@@ 2.2 Forward synthesis
To generate the synthetic observables for this study, we use the DeSIRe code (DeSIRe) to forward-synthesise a set of spectral lines covering atmospheric heights from the deep photosphere to the chromosphere, enabling a comprehensive assessment of model performance at different optical depths.
The simulation outputs are provided on a three-dimensional grid in geometrical height (km). To facilitate forward synthesis with the DeSIRe code, which requires an optical depth scale, we converted the atmospheric stratifications to the required $\log(\tau)$ grid. This was done using the modelador.x utility distributed with DeSIRe, following the methodology outlined in Campbell2021. The final grid has $81$ depth points.
Figure 1: Mean Stokes II profile across the full field of view, averaged over all spatial pixels in the MANCHA simulation snapshot used for training. The vertical dashed lines and labels indicate the most important diagnostic spectral lines. These lines span a broad range of atmospheric heights, from the deep photosphere to the mid-chromosphere, providing constraints for inversion at multiple depths.
@@ 3.2 Transformer architecture
In the transformer framework, the core computation within each attention layer involves three components: queries, keys, and values, as originally introduced by vaswani2017attention. For a given position in the sequence (e.g., a specific wavelength), the query vector determines what the model is looking for, the key vectors determine where that information is located in the input, and the value vectors represent what is retrieved. For a given input sequence $X\in\mathbb{R}^{N_{\lambda}\times d}$ after projection, the queries, keys, and values are computed as
@@
Thus, in summary, our model includes self-attention in the encoder (across wavelengths), cross-attention in the decoder (between depths and spectral features), and self-attention in the decoder (across depths). Each decoder layer performs cross-attention with the common encoder output irrespective of the number of encoder layers, ensuring that all depth queries attend to a shared spectral representation.
@@ 3.4 Training and validation
where $\hat{y}_{i}$ and $y_{i}$ are the predicted and true outputs for the $i$-th training example, and $N$ is the number of examples in the batch. We experimented with both the Adam and AdamW optimizers (Adam). The latter includes a modest weight decay ($1\times 10^{-4}$) to gently penalize overly large weights and improve generalization. We used mixed-precision (16-bit) arithmetic with automatic loss scaling, which speeds up training and reduces memory without sacrificing stability. To prevent any single batch from causing a sudden, oversized update, we cap the combined size of all gradient values at $1.0$ so if the raw gradients exceed this threshold, they’re uniformly scaled down such that the model only makes a controlled adjustment. We performed early stopping based on validation loss, with a maximum of 35 epochs per run.
@@ 4.4 Transformers as a regulariser
4.5 Interpretability of cross-attention
To explore how the transformer achieves its performance, we inspected the cross-attention between optical depth queries and the common encoder output (spectral tokens). In the decoder, each depth query $q_{i}$ is compared with all wavelength keys $k_{j}$ through the scaled dot-product attention mechanism (Eq. 3), producing a weight $a_{ij}$ that expresses how strongly optical depth $i$ attends to wavelength $j$. The resulting attention matrix $A\in\mathbb{R}^{n_{\tau}\times n_{\lambda}}$ can be visualised as a heatmap across wavelength and depth, showing, for each decoder layer and head, how strongly each depth query attends to each wavelength. Here we retrieve the cross-attention from the last decoder layer because it provides the final cross-attention signal; after the layer’s feed-forward sublayer, the resulting decoder states are passed to the output head and thus most directly influence the predictions. Maps can be produced for each head.
We find that each of the four attention heads specialises on different spectral regions and depth ranges. This is illustrated in the sample attention maps shown in Fig. 15 for the Na I D1 line. For visual comparison with response functions, we row-normalise the attention maps across all wavelengths (i.e. $3036$ wavelengths, including those outside the Na I D1 line) per head so that each depth’s most attended wavelength is scaled to unity; this way we can clearly see which wavelength stands out most and how all the others compare to it, without the plot being skewed by differences between depths or heads. For this line, one head concentrates on the wings at photospheric depths (log $\tau\approx-1$ to 2.5-2.5), while another attends primarily to the continuum-forming layers (log $\tau\approx 0$), and while others focus on the line cores at higher layers (up to log $\tau\approx-3.5$). Note that because the maps are row-normalised, they highlight qualitative patterns of focus within each depth and head, rather than giving absolute attention strengths that can be compared across depths or heads. Similar qualitative behaviour is observed for other lines, especially all the upper photospheric lines.
Figure 15: Cross-attention maps for the Na I D1 line retrieved from the final decoder layer, shown separately for each of the four attention heads. Each panel shows the row-normalised attention weight assigned between spectral tokens and depth queries. The maps are computed by averaging over $10000$ Stokes vectors. Each row is scaled by its global maximum; values indicate which wavelengths are preferred at that depth, not how much absolute attention is assigned.
This behaviour is qualitatively consistent with the response functions of these lines; wings are most sensitive to photospheric conditions, while cores are sensitive to higher layers, as demonstrated in Figure 16 for the response function of Stokes $I$ to $T$. This is interpreted as evidence that the transformer assigns weight to spectral regions where the line is known to carry diagnostic information.
Because we jointly predict six atmospheric parameters and attend over per-wavelength tokens that embed all four Stokes components together, this comparison is not one-to-one with parameter-specific response functions. Nevertheless, the alignment between attention maps and known line-formation properties suggests that the transformer has learned to extract information from physically diagnostic parts of the spectrum.
It is important to note that the transformer also contains feed-forward networks in each encoder and decoder layer, so these cross-attention maps should not be over-interpreted as a complete explanation of the model’s learned behaviour. Rather, they are qualitative indicators of which spectral regions and depths are emphasized during the mapping from Stokes profiles to atmospheric parameters.
Figure 16: Response functions of Stokes $I$ to $T$ (upper panel), Stokes $I$ to $v_{\mathrm{LOS}}$ (middle panel), and Stokes $V$ to $B$ (lower panel), for all spectral lines included in the input sequence, computed from a semi-empirical quiet-Sun reference model with $B=500$ G, $v_{\mathrm{LOS}}=0.5$ km/s and $\gamma=45^{\circ}$. The response functions have been normalised by the maximum value for visualisation.
@@ 5 Discussion
In this proof of concept study, we have demonstrated the application of a transformer-based neural network, originally developed for natural language processing, to the inversion problem of inferring stratified atmospheric parameters from full-Stokes, multi-line spectropolarimetric data generated from realistic 3D MHD simulations. Our results show that the transformer model outperforms an MLP baseline in nearly all aspects: reproducing the spatial and vertical morphology of temperature, magnetic field strength, and velocity; achieving higher correlations with the ground truth at all optical depths; and, crucially, delivering physically plausible solutions even in challenging chromospheric layers. The best-performing MLP has approximately $55.2\,\mathrm{M}$ trainable parameters, whereas the winning Transformer has about $5.42\,\mathrm{M}$. Despite being roughly ten times smaller, the Transformer outperforms the MLP. The MLP is parameter-heavy but relatively compute-light, while the Transformer is parameter-light but compute-heavy.
@@
The attention mechanism at the core of the transformer architecture enables the model to leverage non-local spectral features, potentially corresponding to Zeeman splitting or subtle asymmetries, which may be overlooked by models that treat each wavelength as just another input feature, with no explicit bias toward local or long-range spectral relationships. Although a detailed interpretability study is beyond the current scope, initial inspection of the attention maps suggests that the model prioritises physically meaningful regions of the spectrum for inference. This could open new avenues for understanding how neural networks extract physical parameters from complex solar spectra.
@@ 6 Outlook
This study provides the first transformer-based model for full-vector Stokes inversions in solar physics, enabling accurate and efficient inference of atmospheric parameters over many spectral lines and at many optical depths. We anticipate a follow-up study to apply the transformer to real spectropolarimetric observations, such as those from Sunrise III, DKIST, and other current or next-generation facilities. Training on simulations with advanced chromospheric physics will be critical to ensure robust generalisation. An open question is whether transformers might be useful in providing initial guesses for subsequent inversions, though this requires careful testing in future work.
We thank the anonymous referee who helped us significantly improve the manuscript. R.J.C. would like to thank Robert Ryans and Ernst de Mooij for approving use of an NVIDIA Quadro RTX 8000 GPU for model development, as well as technical staff from the Kelvin2 HPC support teams in using their NVIDIA H100 GPUs. R.J.C. thanks Elena Khomenko for permission to use the MANCHA models. This is UKRI-supported research under the Science and Technology Facilities Council (STFC) grants ST/P000304/1 and ST/X000923/1. We acknowledge that the MANCHA simulations used in this paper are produced using the technical expertise and assistance provided by the Spanish Supercomputing Network (Red Española de Supercomputación), as well as the computer resources used: LaPalma Supercomputer, located at the Instituto de Astrofísica de Canarias, and MareNostrum based in Barcelona/Spain. Data availability statement: The code that defined the models generated in this study is available on GitHub. The MANCHA simulation code is available on their public repository. The simulation data used in this study were generated by a third party and are not publicly archived. Access to these data may be available from the original authors on request. {contribution}R.J.C. conceived the project, designed and implemented the machine learning models, conducted the analysis and authored the paper. $M.$ provided critical feedback on the framing of the study and the manuscript. C.Q.N. provided simulation data converted to an optical depth scale that formed the training and test set. All authors approved the manuscript before submission.